OpenStudy (anonymous):

show that x^3 + y^3 >= (x^2)y + (y^2)x

OpenStudy (anonymous):

got any ideas?

OpenStudy (anonymous):

x^2+y^2>=0? then amplify by (x-y)? like a thousand years that i don't do this xd

OpenStudy (anonymous):

x^2y+xy^2-x^3=<y^3

OpenStudy (anonymous):

agreed

OpenStudy (anonymous):

what?

myininaya (myininaya):

that above isn't always true are there any restrictions on x and y?

OpenStudy (anonymous):

well they are intergers

myininaya (myininaya):

that isn't true for all integers you can find a counterexample

myininaya (myininaya):

a counterexample is an example showing the statement is not true for all values within the restriction on the variables

OpenStudy (anonymous):

i think it should be solved using arithmetic or geometric median

OpenStudy (anonymous):

geometric is best I think

myininaya (myininaya):

@zupari It is very easy to think of a counterexample.

OpenStudy (anonymous):

i don't understand

myininaya (myininaya):

try x=-1 and y=0 you will see you expression is not true either the restriction on variables you gave me is wrong or you are expected to find a counterexample

OpenStudy (anonymous):

they are positive intergers

myininaya (myininaya):

that is a different case see if you can use (x+y)^3 in helping you with this problem

OpenStudy (anonymous):

oh wait they are positive real numbers

OpenStudy (anonymous):

tried but no use

OpenStudy (anonymous):

hurry. In 6 hours exam starts

OpenStudy (anonymous):

x>0 and y>0? Then (x-y)^2>=0 its true and x+y>=0 its true So play with 'em

ganeshie8 (ganeshie8):

AM-GM inequality : \[\large x^2 + y^2 \ge 2xy\] \[\large x^2 + y^2-xy \ge xy\] \[\large (x+y)(x^2 + y^2-xy) \ge xy(x+y)\] \[\large x^3 + y^3 \ge x^2y + y^2 x\]

OpenStudy (anonymous):

wow thanks ganeshie8

ganeshie8 (ganeshie8):

notice that the above proof was just messing around the identity : \(\large a^3 + b^3 = (a+b)(a^2-ab+b^2)\)

ganeshie8 (ganeshie8):

also `x+y>=0 ` was used in line 3 the inequality holds only when you multiply both sides by a nonnegative number

OpenStudy (anonymous):

got it